# Charge collection efficiency of thimble ionization chambers exposed to ultra-high dose per pulse

### v1

@@ III.B. Comparison between numerical and analytical model
As highlighted by Fenwick et al.40, when the electron attachment coefficient and the electron mobility are assumed to be independent of the electric field strength and electric field perturbation is neglected, no difference in CCE between positive and negative polarities is expected. The observed polarity effect can be attributed to a variation in the free electron fraction, which depends on the bias voltage polarity. This is a consequence of the dependence of electron transport parameters on the electric field, combined with the presence of electric field perturbations. As a result, CCE is higher when electrons drift toward the central electrode of the chamber. This corresponds to negative sign collected charge, achieved either by applying negative polarity to the external electrode or positive polarity to the collecting electrode.
@@
III.C. Thimble ICs as weighted cylindrical and spherical geometry
@@
In contrast to the PP3D IC, the influence of the selected simulation geometry for the PP IC is less significant. For this chamber, the maximum discrepancy between the volume-weighted average and the simplified 2D geometry is 0.2 % for CCE and 0.2 % for . The maximum difference between the simplified 2D geometry and the complete 2D geometry reaches 1.7 % for CCE and 1.9 % for . For both ICs, the presence of a guard ring leads to a reduction in both CCE and .
Although explicitly indicated, the following sections use the complete 2D geometry for the numerical simulations.
III.D. Comparison between experimental and simulated charge collection efficiency and polarity effect
The experimental and simulated CCE and are shown in Figure 4 for the PP3D IC (SN 152986) and the PP IC (SN 170313) as a function of the DPP. These simulations were performed using Boissonnat et al.38 mobilities together with a volume recombination coefficient of 1.410-12 m3 s-1. As expected, the CCE is larger when the collected charge sign is negative, resulting in . On average, discrepancies between experimental and simulated CCE () are 1.4 % (7.0 %) and 1.6 % (3.2 %) for the PP3D and PP ICs, respectively.
@@ III.D. Comparison between experimental and simulated charge collection efficiency and polarity effect
A similar agreement is found when using the ion mobilities reported by Zhang et al.37 together with an ion-ion recombination coefficient of 0.810-12 m3 s-1. On average, discrepancies between experimental and simulated CCE () are 2.0 % (7.3 %) and 1.7 % (3.6 %) for the PP3D and PP ICs, respectively.
Regarding pulse duration, CCE increases as pulse duration becomes longer. Figure 5 shows the influence of pulse duration on both CCE and for the two ICs under a bias voltage of 500 V. In the overlapping DPP range (1.7 Gy to 3.1 Gy per pulse) differences in CCE up to 4.8 % and 4.2 % were found between 1.0 s and 1.9 s for the PP3D and PP ICs, respectively. The simulation successfully reproduces the observed trend and the relative variation in CCE due to pulse duration with discrepancies below 2 % at 500 V.
The experimental difference in CCE () on average for DPP from 1.6 Gy to 6 Gy between the two samples of each chamber type at 500 V was 4.5 % (2.8 %) and 1.4 % (1.6 %) for the PP3D and the PP ICs, respectively. This intra-model variability aligns with previous findings by Bourgouin et al.2 who reported variations up to 20 % at 6 Gy per pulse in parallel-plate ICs.
@@
III.E. Time-resolved currents
@@
Similar to the behavior observed in parallel-plate ICs 4, the time-resolved current of thimble ICs can be separated into two distinct contributions. During the pulse, a large current is generated by the rapid drift of free electrons within the IC. Variations in the pulse delivery time profile on the nanosecond scale significantly affect the IC’s instantaneous current due to the high mobility of electrons. After the beam pulse ends, a prolonged ion tail is observed, resulting from the slower drift of ion species. Figure 6 shows the time-resolved currents for the PP3D (panel A) and the PP IC (panel B) for +300 V and -300 V bias voltage irradiated with a DPP of 3 Gy and a pulse duration of 1.9 s. Both chambers exhibit higher instantaneous current during the pulse at negative voltage. This corresponds to electrons drifting toward the central electrode, where the electric field is stronger. Additionally, charge collection is faster in this configuration, due to the combined effects of electric field perturbation and the slower drift of positive ions toward regions with lower field strength. When comparing both chambers under the same bias voltage, the PP chamber shows a faster charge collection time than the PP3D, primarily due to its smaller external radius.
III.F. Invariant rules for the charge collection efficiency of ionization chambers
@@
Previous work of Kranzer et al.5 showed that the CCE of two parallel-plate ICs and with distance between electrodes and biased with voltages and are approximately equal under same irradiation conditions. This was verified both experimentally and with numerical simulation, giving a scaling rule as:
@@
Bourgouin et al. 2 evaluated this relationship across a broader set of parallel-plate ICs, demonstrating reasonable agreement among various ICs models. This invariant rule seems to be connected to the Boag model, where the CCE is a function of the dimensionless parameter (Equation 7), preserving this functional dependency. However, the applicability of the Boag model is restricted to DPPs values where the electric field perturbation can be neglected. Furthermore, even in the low DPP regime, its validity for air-vented ICs is compromised as it disregards the contribution of free electrons that drift across the IC without undergoing attachment 42.
@@
Surprisingly, this simple invariant rule seems to work for air-vented parallel-plate ICs at large DPPs, where electric field perturbation largely impacts the charge transport. Recently, Fenwick et al. 40 have shown that the CCE of a cylindrical IC can be related to the CCE of a parallel-plate IC at the same bias voltage if the geometrical parameters and are the same when the electric field perturbation is negligible.
@@
Figure 7 displays the CCE measurements of several samples of Advanced Markus parallel-plate ICs (nominal distance between electrodes of 1 mm) from Bourgouin et al.2 and two 1 mm parallel-plate prototypes from Kranzer et al.5, using 250 V and 500 V bias voltages. The CCE of the PP3D operated at 400 V ( mm2/kV) closely follows the trend observed for the parallel-plate ICs at 250 V (mm2/kV). Similarly, the PP IC operated at 300 V (mm2/kV) shows a similar trend to that of the parallel-plate ICs at 500 V (mm2/kV). For the PP3D, a slight deviation at large DPPs values might be attributed to the influence of the guard ring and the spherical tip on the CCE. In fact, the simulation using the 1D geometry yields a better agreement with the CCE of the parallel-plate ICs at 250 V. Here, the 1D simulation was performed using the real IC volume, in order to obtain a similar released charge in the chamber for a given DPP as in the parallel-plate ICs.
IV. Discussion
@@
According to the results presented in section III.C., the CCE of a thimble IC can be reasonably approximated by the volume-weighted average CCE of simple spherical and cylindrical geometries if no guard ring is present. The discrepancy between the weighted approach and the full simulation is surprisingly smaller than 3 % whenever the guard ring does not substantially affect the chamber response. The presence of a guard ring electrode introduces an insensitive region within the chamber (dead volume) and affects charge transport, leading to a reduction in both polarity correction factor and CCE. Therefore, in order to accurately reproduce the CCE of thimble ICs, especially when the dead volume is an important fraction of the total volume (nearly 30 % for the PP3D compared to 6 % for the PP), it is advisable to include the guard ring in the simulation. This dead volume also explains the small difference observed in the calibration coefficients between the two ICs types of about +4.6 % experimentally and +5.0 % predicted by simulation, while the physical volume difference between the ICs is about -15 %.
@@
As shown in Section III.D., both experimental data and simulations indicate that the studied ICs exhibit an increasing polarity effect correction factor, , with increasing DPP. The numerical model attributes a large fraction of this polarity effect to volume recombination processes that depend on the polarity of the applied bias voltage, contrary to low DPP where volume recombination no longer plays a major role and the polarity effect is primarily caused by extracavitary charge contributions. Consequently, for thimble-type chambers operating at high DPP, there is an interplay between the CCE and the polarity correction factor, making their separation and interpretation more complex. The discrepancies observed in terms of might indicate limitations in the implementation of the modeled IC geometry. However, as reported previously by Bourgouin et al.2, parallel-plate ICs irradiated with large DPPs (sufficient to cause a reduction in CCE) also exhibit an increasing polarity effect with DPP. In addition to polarity dependent charge transport and recombination, other factors may contribute to the observed polarity effect. Further research into the underlying mechanisms of polarity effect in ICs exposed to ultra-high-DPP conditions is recommended.
@@
The invariant rules used to relate the CCE of different ICs, as discussed in Section III.F., were previously identified by Kranzer et al.5 for parallel-plate ICs. The use of the parameter , with units of length squared over bias voltage, appears to provide an invariant for CCE, with a rough agreement of approximately 5 %. Extending this approach to cylindrical chambers has demonstrated its usefulness not only for comparing different thimble-type ICs, but also for establishing relationships between parallel-plate and thimble chambers. However, since the parameter is derived from an idealized cylindrical geometry, the resulting scaling rule should be expected to provide only an approximate prediction for chamber intercomparison.
@@ Acknowledgments
The data supporting this findings is available at the following repository: https://doi.org/10.5281/zenodo.17737817

### v2

@@ 3.2 Comparison between numerical and analytical model
As highlighted by Fenwick et al. [40], when the electron attachment coefficient and the electron mobility are assumed to be independent of the electric field strength and electric field perturbation is neglected, no difference in CCE between positive and negative polarities is expected. The observed polarity effect can be attributed to a variation in the free electron fraction, which depends on the bias voltage polarity. This is a consequence of the dependence of electron transport parameters on the electric field, combined with the presence of electric field perturbation. As a result, CCE is higher when electrons drift toward the central electrode of the chamber. This corresponds to negative sign collected charge, achieved either by applying negative polarity to the external electrode or positive polarity to the collecting electrode.
@@
3.3 Polarity effect in cylindrical ionization chambers
For a better comprehension of the polarity effect arising from recombination in cylindrical ICs, charge transport was studied in a 1D cylindrical IC with geometrical characteristics similar to those of the PP3D IC. Numerical simulations were performed using a DPP of 3 Gy, a pulse duration of 1.9 s and a bias voltage of 300 V for both positive and negative polarity. These results were obtained with the transport parameters listed in Table 3.
It is important to emphasize that the dominant contribution to charge loss comes from ion-ion recombination, as electron–ion recombination is a subdominant process. Ion-ion recombination occurs through the overlap of positive and negative ion densities within the active volume of the chamber, quantified by whose time evolution is shown in Figure 3 for both polarities. The overlap between the negative and positive ion distributions reaches its maximum value at the end of the radiation pulse. As clearly illustrated in Figure 3 the ion density overlap –and therefore the recombination effect– is greater when the collected charge at the central electrode is positive.
The simulated charge carrier densities and electric field as a function of the radial coordinate are shown in Figure 4 at the end of the radiation pulse. When the collected charge is negative (panel A), the main region of ion overlap is located closer to the central electrode, whereas it shifts toward the external electrode (panel B) when the collected charge is positive. In Figure 4 the electric field perturbation caused by net charge imbalance leads to a significantly reduced field value in the ion overlap region after the pulse delivery. Nevertheless, the higher electric field magnitude in the overlap region for negative polarity (panel A) has a twofold effect: it produces a faster drift of the charges –thereby reducing both the ion-overlap time and the electron transit time– and it increases the free electron fraction, which again reduces charge loss by depleting the negative ion population. This second effect is enhanced by the increase in electron lifetime in air with electric field magnitude at intermediate values (0.1 kV/mm to 1.0 kV/mm), resulting in a lower amount of negative ions when electrons drift toward the central electrode. These effects contribute to a higher recombination in a thimble chamber when the polarity of the collected charge is positive. The same physical principles apply for lower DPP deliveries, leading to a higher CCE that still retains a polarity dependence as shown in Figure 5.
3.4 Description of thimble ICs as weighted cylindrical and spherical geometry
@@
In contrast to the PP3D IC, the influence of the selected simulation geometry for the PP IC is less significant. For this chamber, the maximum discrepancy between the volume-weighted average and the simplified 2D geometry is 0.2 % for CCE and 0.2 % for . The maximum difference between the simplified 2D geometry and the complete 2D geometry reaches 1.7 % for CCE and 1.9 % for . For both ICs, the presence of a guard ring leads to a reduction in both CCE and . Unless explicitly stated, the following sections use the complete 2D geometry for the numerical simulations.
3.5 Comparison between experimental and simulated charge collection efficiency and polarity effect
The experimental and simulated CCE and are shown in Figure 6 for the PP3D IC (SN 152986) and the PP IC (SN 170313) as a function of the DPP. These simulations were performed using Boissonnat et al. [38] mobilities together with a volume recombination coefficient of 1.410-12 m3 s-1. As expected, the CCE is larger when the collected charge sign is negative, resulting in . On average, discrepancies between experimental and simulated CCE () are 1.4 % (7.0 %) and 1.6 % (3.2 %) for the PP3D and PP ICs, respectively.
@@ 3.5 Comparison between experimental and simulated charge collection efficiency and polarity effect
A similar agreement is found when using the ion mobilities reported by Zhang et al. [37] together with an ion-ion recombination coefficient of 0.810-12 m3 s-1. On average, discrepancies between experimental and simulated CCE () are 2.0 % (7.3 %) and 1.7 % (3.6 %) for the PP3D and PP ICs, respectively.
Regarding pulse duration, CCE increases as pulse duration becomes longer for a given DPP. Figure 7 shows the influence of pulse duration on both CCE and for the two ICs under a bias voltage of 500 V. In the overlapping DPP range (1.7 Gy to 3.1 Gy per pulse) differences in CCE up to 4.8 % and 4.2 % were found between 1.0 s and 1.9 s for the PP3D and PP ICs, respectively. The simulation successfully reproduces the observed trend and the relative variation in CCE due to pulse duration with discrepancies below 2 % at 500 V.
The experimental difference in CCE () on average for DPP from 1.6 Gy to 6 Gy between the two samples of each chamber type at 500 V was 4.5 % (2.8 %) and 1.4 % (1.6 %) for the PP3D and the PP ICs, respectively. This intra-model variability aligns with previous findings by Bourgouin et al. [2] who reported variations up to 20 % at 6 Gy per pulse in parallel-plate ICs.
@@
3.6 Time-resolved currents
@@
Similar to the behavior observed in parallel-plate ICs [4], the time-resolved current of thimble ICs can be separated into two distinct contributions. During the pulse, a large current is generated by the rapid drift of free electrons within the IC. Variations in the pulse delivery time profile on the nanosecond scale significantly affect the IC’s instantaneous current due to the high mobility of electrons. After the beam pulse ends, a prolonged ion tail is observed, resulting from the slower drift of ion species. Figure 8 shows the time-resolved currents for the PP3D (panel A) and the PP IC (panel B) for +300 V and -300 V bias voltage irradiated with a DPP of 3 Gy and a pulse duration of 1.9 s. Both chambers exhibit higher instantaneous current during the pulse at negative voltage. This corresponds to electrons drifting toward the central electrode, where the electric field is stronger. Additionally, charge collection is faster in this configuration, due to the combined effects of electric field perturbation and the slower drift of positive ions toward regions with lower field strength. When comparing both chambers under the same bias voltage, the PP chamber shows a faster charge collection time than the PP3D, primarily due to its smaller external radius.
3.7 Invariant rules for the charge collection efficiency of ionization chambers
@@
Previous work of Kranzer et al. [5] showed that the CCE of two parallel-plate ICs and with distance between electrodes and biased with voltages and are approximately equal under same irradiation conditions. This was verified both experimentally and with numerical simulation, giving a scaling rule as:
@@
Bourgouin et al. [2] evaluated this relationship across a broader set of parallel-plate ICs, demonstrating reasonable agreement among various ICs models. This invariant rule seems to be connected to the Boag model, where the CCE is a function of the dimensionless parameter (Equation 7), preserving this functional dependency. However, the applicability of the Boag model is restricted to DPPs values where the electric field perturbation can be neglected. Furthermore, even in the low DPP regime, its validity for air-vented ICs is compromised as it disregards the contribution of free electrons that drift across the IC without undergoing attachment [42].
@@
Surprisingly, this simple invariant rule seems to work for air-vented parallel-plate ICs at large DPPs, where electric field perturbation largely impacts the charge transport. Recently, Fenwick et al. [40] have shown that the CCE of a cylindrical IC can be related to the CCE of a parallel-plate IC at the same bias voltage if the geometrical parameters and are the same when the electric field perturbation is negligible.
@@
Figure 9 displays the CCE measurements of several samples of Advanced Markus parallel-plate ICs (nominal distance between electrodes of 1 mm) from Bourgouin et al. [2] and two 1 mm parallel-plate prototypes from Kranzer et al. [5], using 250 V and 500 V bias voltages. The CCE of the PP3D operated at 400 V ( mm2/kV) closely follows the trend observed for the parallel-plate ICs at 250 V (mm2/kV). Similarly, the PP IC operated at 300 V (mm2/kV) shows a similar trend to that of the parallel-plate ICs at 500 V (mm2/kV). For the PP3D, a slight deviation at large DPPs values might be attributed to the influence of the guard ring and the spherical tip on the CCE. In fact, the simulation using the 1D geometry yields a better agreement with the CCE of the parallel-plate ICs at 250 V. Here, the 1D simulation was performed using the real IC volume, in order to obtain a similar released charge in the chamber for a given DPP as in the parallel-plate ICs.
4 Discussion
@@
According to the results presented in section 3.4, the CCE of a thimble IC can be reasonably approximated by the volume-weighted average CCE of simple spherical and cylindrical geometries if no guard ring is present. The discrepancy between the weighted approach and the full simulation is surprisingly smaller than 3 % whenever the guard ring does not substantially affect the chamber response. The presence of a guard ring electrode introduces an insensitive region within the chamber (dead volume) and affects charge transport, leading to a reduction in both polarity correction factor and CCE. Therefore, in order to accurately reproduce the CCE of thimble ICs, especially when the dead volume is an important fraction of the total volume (nearly 30 % for the PP3D compared to 6 % for the PP), it is necessary to include the guard ring in the simulation. This dead volume also explains the small difference observed in the calibration coefficients between the two ICs types of about +4.6 % experimentally and +5.0 % predicted by simulation, while the physical volume difference between the ICs is about -15 %.
@@
As shown in Section 3.5, both experimental data and simulations indicate that the studied ICs exhibit an increasing polarity effect correction factor, , with increasing DPP. The numerical model attributes a large fraction of this polarity effect to volume recombination processes that depend on the polarity of the applied bias voltage, contrary to low DPP where volume recombination no longer plays a major role and the polarity effect is primarily caused by extracavitary charge contributions. Consequently, for thimble-type chambers operating at high DPP, there is an interplay between the CCE and the polarity correction factor, making their separation and interpretation more complex.
The discrepancies observed in terms of might indicate limitations
in the implementation of the modeled IC geometry. However, as reported previously by Bourgouin et al. [2], parallel-plate ICs irradiated with large DPPs also exhibit an increasing polarity effect with DPP. In addition to polarity dependent charge transport and recombination, other factors may contribute to the observed polarity effect. In order to understand the mechanisms leading to the polarity effect observed in IC exposed to ultra-high-DPP further research is required.
@@
The invariant rules used to relate the CCE of different ICs, as discussed in Section 3.7, were previously identified by Kranzer et al. [5] for parallel-plate ICs. The use of the parameter , with units of length squared over bias voltage, appears to provide an invariant for CCE, with a rough agreement of approximately 5 %. Extending this approach to cylindrical chambers has demonstrated its usefulness not only for comparing different thimble-type ICs, but also for establishing relationships between parallel-plate and thimble chambers. However, since the parameter is derived from an idealized cylindrical geometry, the resulting scaling rule should be expected to provide only an approximate prediction for chamber intercomparison.
@@ Acknowledgments
The data supporting these findings is available at the following repository: https://doi.org/10.5281/zenodo.17737817. The numerical code is available at https://github.com/JPazMartin/ICSimFEM.
Conflict of Interest
The authors have no relevant conflicts of interest to disclose.